Mass of a nucleus of an atom is 3.87 x 10^-23. This nucleus contains 10 neutrons. Identify the element.


1
Expert's answer
2022-03-22T17:01:05-0400

Mass of nucleus = "3.87 \times 10^{-23} g"


Mass of one neutron = "1.67 \times 10 ^{-24}" g


Total mass of 10 neutron = "1.67 \times 10 ^{-23}" g


Mass of nucleus = Z(mass of one proton) + (A-Z) mass of one neutron


"3.87 \times 10 ^{-23}= (Z\times 1.67 \times 10^{-24} )+ 1.67 \times 10 ^{-23}"


"Z =\frac{ 2.2\times 10 ^{-23}}{1.67 \times 10^-24}= 13"


So number of proton = 13

So , element is Aluminium.